Find the length of the major axis.
(x - 8)^2/49+(y+6)^2/121=1

Answer by stanbon(75887)

Find the length of the major axis.
(x - 8)^2/49+(y+6)^2/121=1
semi-major axis = sqrt(121) = 11
-----
Ans:: length of major axis = 2*11 = 22
----------------
Cheers,
Stan H.
---------------